OpenStudy (anonymous):

What is /27 rounded to the nearest tenth?Select one of the options below as your answer:A. 5.1B. 5.2C. 5.3D. 5.4

OpenStudy (anonymous):

?/27

OpenStudy (anonymous):

it supposed to be a divsion sign , sorry .

hero (hero):

It's Michael Jordan, the greatest of all time :D

OpenStudy (anonymous):

lololol

OpenStudy (anonymous):

\[\sqrt{27} thats how its suppose \to look at Kreshnik .\]

OpenStudy (anonymous):

thats the whole question .

hero (hero):

Okay, so break out your calc. The best you could do before breaking out the calc is this: \[\sqrt{27} = \sqrt{9 \times 3} = \sqrt{9} \dot\ \sqrt{3} = 3\sqrt{3}\]

OpenStudy (anonymous):

so what would the answer be o.o ?

hero (hero):

Where is your calc? You have a computer so I know you have one.

OpenStudy (anonymous):

the answer is D :)

hero (hero):

Hint: It's not D

OpenStudy (anonymous):

idk how to do it /:

OpenStudy (anonymous):

i found the calculator its 5.196 so the answer would be B. :)